OpenStudy (anonymous):

how many different 5-person basketball teams can a coach choose from among 7 players?

OpenStudy (solomonzelman):

7 C 5

OpenStudy (solomonzelman):

(combinations)

OpenStudy (anonymous):

do I put that in a calculator?

OpenStudy (solomonzelman):

Well, I would just do, \(\LARGE\color{blue}{ 7~C~5= \frac{7!}{5!(7-5)!} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ 7~C~5= \frac{7!}{5!(7-5)!}=\frac{7!}{5!\times2!} =\frac{5!\times 6 \times 7}{5!\times2!} }\)

OpenStudy (solomonzelman):

Can you finish it ?

OpenStudy (anonymous):

yes thanks so much!

OpenStudy (solomonzelman):

Anytime !